STUDIES OF VIBRATIONAL EXCITATIONS OF CH_{5}^{+} BY DIFFUSION MONTE CARLO TECHNIQUES

\underline{CHARLOTTE E. HINKLE}, ANNE B. McCOY, Department of Chemistry, The Ohio State University, Columbus, OH 43210.

Understanding the vibrational spectrum of CH_{5}^{+} is a question of great interest. There are currently only a few spectra of CH_{5}^{+}. These include a high resolution spectrum observed by Oka^{a}, a high resolution, jet cooled spectrum that is only partially assigned, and the broad peaks in the low resolution spectrum that is assigned.^{b} The method that we employ for obtaining excited state structures of CH_{5}^{+} is Diffusion Monte Carlo (DMC). DMC has been effectively applied to finding structural information about the ground state of CH_{5}^{+}, and we are currently using extensions of this technique to probe CH excited states of CH_{5}^{+}. Of particular interest is the broad CH stretching band observed between 2200 and approximately 3200 cm^{−1}, and a broad bending region centered at 1200 cm^{−1}.^{c} We will study the nature of states that have a single node along a single coordinate such as the CH stretching and bending motions.

aE. T. White, J. Tang, T. Oka, Science 284, 135 (1999).

bX. Huang., A. B. McCoy, J. M. Bowman, L. M. Johnson, C. Savage, F. Dong, D. J. Nesbitt, Science 311, 60 (2006).

cO. Asvany, P. Kumar P, B. Redlich, I. Hesemann, S. Schlemmer, D. Marx, Science 309, 1219 (2005).